# Stable generative modeling using Schrödinger bridges

### v2

@@ 3. Approximating the conditional mean
In order to implement (4), we need to define $m_{\epsilon}(x)$ and $\Sigma(x)$ for $x\in\mathbb{R}^{d}$. In this section, we discuss how one can obtain these functions from the training samples $\{x^{(i)}\}_{i=1}^{M}$ and the Markov chain approximation (8).
We introduce the vector $t_{\epsilon}(x)\in\mathbb{R}^{M}$ with entries
(9) $t_{\epsilon,i}(x)=\exp\left(-\frac{1}{2\epsilon}(x^{(i)}-x)^{\top}\left(K(x)+K(x^{(i)})\right)^{-1}(x^{(i)}-x)\right)$
for $i=1,\ldots,M$. We then define the probability vector using the Sinkhorn weights, $v_{\epsilon}$, obtained in (8), i.e.,
(10) $p_{\epsilon}(x)=\frac{D(v_{\epsilon})t_{\epsilon}(x)}{v_{\epsilon}^{\top}t_{\epsilon}(x)}\in\mathbb{R}^{M}$
for all $x\in\mathbb{R}^{d}$. This vector gives the transition probabilities from $x$ to $\{x^{(i)}\}_{i=1}^{M}$ and provides a finite-dimensional approximation to the conditional probability distribution $\pi_{\epsilon}(\cdot|x)$ of the true underlying diffusion process; i.e., the semigroup $\exp(\epsilon\mathcal{L})$ with generator $\mathcal{L}$ corresponding to the generalized reversible diffusion process (6). See Section 4 for more details.
Finally, using the probability vector $p_{\epsilon}$ and introducing the data matrix of samples
@@
The construction of the conditional mean $m_{\epsilon}(x)$ is known as the barycentric projection of the entropy-optimally coupling [48, 42]. In optimal transport, $v_{\epsilon}$ plays the role of the optimizer of the dual problem.
@@ 3.1. Sampling algorithms
We now present our main Langevin sampling strategies based on the previously introduced probability vector $p_{\epsilon}(x)$ in (10). Sampling schemes of the form (4) have the same drift term $m_{\epsilon}(x)$, but differ in the way the diffusion matrix $\Sigma(x)$ is defined. We consider a data-unaware diffusion as well as a data-aware diffusion which turns out to be advantageous in generating new samples from the data distribution $\pi$. See the numerical experiments in Section 6.
@@ 3.1.1. Langevin sampler with data-unaware diffusion
(13) $X_{n+1}=X_{n}+\Delta\tau\left(\frac{m_{\epsilon}(X_{n})-X_{n}}{\epsilon}\right)+\sqrt{2K(X_{n})}\Xi_{n}$
@@
(14) $s_{M}(x)=\frac{m_{\epsilon}(x)-x}{\epsilon}$
@@
Let us briefly discuss the qualitative behavior of the time-stepping method (15) as a function of $\epsilon>0$. For large $\epsilon$, the expected value $m_{\epsilon}(x)$ will become essentially independent of the current state $X_{n}$ and the diffusion process will sample from a centred Gaussian. For $\epsilon\to 0$, on the other hand, the probability vector $p_{\epsilon}(x)$ can potentially degenerate into a vector with a single entry approaching one with all other entries essentially becoming zero. Hence a key algorithmic challenge is to find a good value for $\epsilon$ and a suitable $K(x)$, which guarantee both good mixing and accuracy, i.e., $X_{n}\sim\pi$ as $n\to\infty$.
@@ 3.1.2. Langevin sampler with data-aware diffusion
(16) $C(x)=\epsilon^{-1}(\mathcal{X}-m_{\epsilon}(x)1_{M}^{\top})D(p_{\epsilon}(x))(\mathcal{X}-m_{\epsilon}(x)1_{M}^{\top})^{\top}\in\mathbb{R}^{d\times d},$
which is the (scaled) covariance matrix associated with the probability vector $p_{\epsilon}(x)$. Here $1_{M}\in\mathbb{R}^{M}$ denotes the $M$-dimensional vector of ones. Therefore, one can more directly implement a Gaussian approximation associated with the transition probabilities $p_{\epsilon}(x)$ and introduce the update
(17) $X_{n+1}=X_{n}+\Delta\tau\left(\frac{m_{\epsilon}(X_{n})-X_{n}}{\epsilon}\right)+\sqrt{C(X_{n})}\Xi_{n}.$
@@
(18) $X_{n+1}=m_{\epsilon}(X_{n})+\sqrt{C(X_{n})}\Xi_{n},$
which we found to work rather well in our numerical experiments since it directly captures the uncertainty contained in the data-driven coupling $P_{\epsilon}$. The scheme (18) corresponds to setting $\Sigma(x)=C(x)$ in (4). Also note that the scheme (18) still depends on $K(x)$ through the probability vector $p_{\epsilon}(x)$.
@@ 3.2. Algorithmic properties
We briefly discuss several important properties on the stability and the ergodicity of the proposed Langevin samplers.
The following Lemma establishes that, since each $p_{\epsilon}(x)$ is a probability vector, $m_{\epsilon}(x)=\mathcal{X}p_{\epsilon}(x)$ is a convex combination of the training sample $\{x^{(i)}\}_{i=1}^{M}$.
@@
The Lemma follows from the definition (12), which we write as $m_{\epsilon}(x)=\sum_{j=1}^{M}x^{(j)}p_{\epsilon_{j}}(x)$, and the fact that $p_{\epsilon}(x)$ is a probability vector with $0\leq p_{\epsilon_{j}}(x)\leq 1$ for all $\epsilon>0$ and all $x\in\mathbb{R}^{d}$. ∎
@@
Let us assume that the data generating density $\pi$ has compact support. Then the time-stepping method (15) possesses a unique invariant measure and is geometrically ergodic provided the norm of the symmetric positive matrix $K(x)$ is bounded from above and below for all $x\in\mathbb{R}^{d}$.
@@
Consider the Lyapunov function $V(x)=1+\|x\|^{2}$ and introduce the set
$\mathcal{C}=\{x:\|x\|\leq R\}$
for suitable $R>0$. Since $m_{\epsilon}(X_{n})\in\mathcal{C}_{M}$ and $\pi$ has compact support, one can find a radius $R>0$, which is independent of the training data $\{x^{(i)}\}_{i=1}^{M}$, such that $\mathcal{C}_{M}\subset\mathcal{C}$ and
$\mathbb{E}[V(X_{n+1})|X_{n}]\leq\lambda V(X_{n})$
for all $X_{n}\notin\mathcal{C}$ with $10\leq\lambda<$ 1. Furthermore, because of the additive Gaussian noise in (15), there is a probability density function $\nu(x)$ and a constant $\delta>0$ such that
${\rm n}(x^{\prime};m_{\epsilon}(x),2\epsilon K(x))\geq\delta\nu(x^{\prime})$
for all $x,x^{\prime}\in\mathcal{C}$. Here ${\rm n}(x;m,\Sigma)$ denotes the Gaussian probability density function with mean $m$ and covariance matrix $\Sigma$. In other words, $\mathcal{C}$ is a small set in the sense of [38]. Geometric ergodicity follows from Theorem 15.0.1 in [38]. See also the self-contained presentation in [36]. ∎
@@
(21b) $X_{n+1}$ $=\mathcal{X}p_{\epsilon}(X_{n+1/2}).$
These split-step schemes have been used in our numerical experiments.
@@ 3.3. Variable bandwidth diffusion
It is an active area of research to select a $\rho$ that increases the spectral gap of $\mathcal{L}$ while not increasing computational complexity. A larger spectral gap implies a faster convergence rate [43], indicating that the generated samples are closer to the reference at a finite time, exhibiting a high accuracy. We demonstrate numerically in Section 6 that $\rho$ can indeed be used to increase the sampling accuracy. More specifically, the bandwidth $\rho(x)$ is chosen as
@@
$\tilde{\pi}_{i}=Z^{-1}\pi_{i},\qquad Z:=\frac{1}{M}\sum_{j}\pi_{j}.$
@@
$\rho_{i}=\tilde{\pi}_{i}^{\beta}=Z^{-\beta}\pi_{i}^{\beta}$
@@
4. Connections to the generator of Langevin dynamics
In this section, we discuss the proposed methodology in the limit $M\to\infty$ under the idealised assumption that the target distribution $\pi$ is in fact known. It is well-known that $K=I$ implies that the Schrödinger bridge approximates the semi-group corresponding to the generator of standard Langevin dynamics [34, 54]. Employing a variable bandwidth (22) instead alters the semi-group and thus the underlying generator. However, since in both cases the density $\pi(x)$ remains invariant, this is not an issue as we are only concerned with drawing samples from the distribution and not in reproducing any dynamical features.
More precisely, using (22) in (7), we consider the associated Markov chain approximation given by (8). We formally take the limit $M\to\infty$ and denote the limiting operator by $\mathcal{P}_{\epsilon}$ [54]. One formally obtains
$\mathcal{P}_{\epsilon}=e^{\epsilon\mathcal{L}},$
where $\mathcal{L}$ denotes the generator defined by
(25) $\mathcal{L}f=\pi^{-1}\nabla\cdot(\pi\rho\nabla f)=\nabla\cdot(\rho\nabla f)+\rho\nabla\log\pi\cdot\nabla f.$
We note that $\mathcal{L}$ is self-adjoint (reversible) with respect to the $\pi$-weighted inner product. It is more revealing to consider the dual operator
$\mathcal{L}^{\dagger}\mu=\nabla\cdot(\mu\rho\nabla(\log\mu-\log\pi))$
and the associated mean-field evolution equation
$\dot{X}_{\tau}=\rho(X_{\tau})\left(\nabla(\log\pi(X_{\tau})-\log\mu(X_{\tau})\right),\qquad X_{0}\sim\mu_{0},$
which implies the invariance of $\mu=\pi$.
Please also note that the generator (25) corresponds to the diffusion process (6) with $K(x)=\rho(x)I$ in Itô form and to
$\dot{X}_{\tau}=\rho(X_{\tau})\nabla\log\pi(X_{\tau})+\sqrt{2\rho(X_{\tau})}\diamond\dot{W}_{\tau}$
in Klimontovitch form [23]. The drift term in the Itô formulation (6) can be expressed as
@@ 4. Connections to the generator of Langevin dynamics
and, hence, (14) provides a finite $M$ approximation to the score function
$s(x)=\mathcal{L}\,\mbox{Id}(x).$
However, recall that the recursive Langevin sampler (15) relies on a direct time-stepping approach to (6) for $K(x)=\rho(x)I$ and is based on
$m_{\epsilon}(x)=\mathcal{X}p_{\epsilon}(x)\approx\exp(\epsilon\mathcal{L})\,\mbox{Id}(x),$
instead of an approximation of the drift term on the right hand side of (6) via (26) followed by a Euler–Maruyama time discretization. However, both approaches are formally consistent in the limit $\epsilon\to 0$ as
$\exp(\epsilon\mathcal{L})\,\mbox{Id}(x)\approx x+\epsilon s(x).$
While we have focused on the particular choice $K(x)=\rho(x)I$ in this section for simplicity, the discussion naturally extends to general symmetric positive definite matrices $K(x)$.
@@ 5. Bayesian inference and conditional sampling
where $V(x)$ denotes a known potential such as the negative log-likelihood function in case of Bayesian inference with $\pi$ taking the role of the prior distribution. We obtain, for example, the adjusted time-stepping scheme
$X_{n+1/2}$ $=X_{n}-\epsilon\nabla V(X_{n})+\sqrt{2}\,\Xi_{n},$
$X_{n+1}$ $=\mathcal{X}p_{\epsilon}(X_{n+1/2})$
@@
$x^{\ast}=\arg\min_{x}\left\{V(x)-\log\pi(x)\right\}$
for given potential $V(x)$ and regulariser $-\log\pi(x)$. Again assuming that only samples $\{x^{(i)}\}_{i=1}^{M}$ of $\pi$ are available, an algorithm for approximating $x^{\ast}$ can be defined as follows:
$x_{n+1/2}$ $=x_{n}-\epsilon\nabla V(x_{n}),$
$x_{n+1}$ $=\mathcal{X}p_{\epsilon}(x_{n+1/2}).$
Here $m_{\epsilon}(x)=\mathcal{X}p_{\epsilon}(x)$ approximates the optimization update associated with the regulariser $-\log\pi(x)$. The iteration is stable and any limiting point $x_{\infty}$ is contained in the convex hull of the data points $\{x^{(i)}\}_{i=1}^{M}$.
Alternatively, the sampling schemes (20) and (21) can also easily be extended to conditional generative modeling. More specifically, consider a random variable in $x=(y,z)$, which we condition on the second component for given $z=z^{\ast}$. In other words, we wish to sample from $\pi(y|z^{\ast})$ given samples $x^{(i)}=(y^{(i)},z^{(i)})$, $i=1,\ldots,M$, from the joint distribution $\pi(x)=\pi(y,z)$.
In order to perform the required conditional sampling, we propose a method which combines approximate Bayesian computation (ABC) with our diffusion map based sampling algorithm. Let us assume that we can generate $M$ samples $x^{(i)}=(y^{(i)},z^{(i)})$, $i=1,\ldots,M$, from a joint distribution $\pi(y,z)$, which we then wish to condition on a fixed $z^{\ast}$. As before, we construct vectors of transition probabilities $p_{\epsilon}(x)\in\mathbb{R}^{M}$ based on the samples $\{x^{(i)}\}_{i=1}^{M}$. We assume that the bandwidth parameter $\epsilon$ used in the diffusion map approximation is also applied in the ABC misfit function, that is,
$L(z,z^{\ast})=\frac{1}{2\epsilon}\|z-z^{\ast}\|^{2}.$
This suggests the following split-step approximation scheme. Given the last sample $X_{n}=(Y_{n},Z_{n})$, we first update the $z$-component using
$\hat{Z}_{n}=Z_{n}-\epsilon\nabla_{z}L(Z_{n},z^{\ast})=z^{\ast}.$
In other words, we replace the current $X_{n}$ by $\hat{X}_{n}=(Y_{n},z^{\ast})$. Next we apply the split-step scheme (20) to $\hat{X}_{n}$, i.e.,
$X_{n+1/2}$ $=\hat{X}_{n}+\sqrt{2K(\hat{X}_{n})}\Xi_{n},$
$X_{n+1}$ $=\mathcal{X}p_{\epsilon}(X_{n+1/2}),$
where we have again $\mathcal{X}=(x^{(1)},\ldots,x^{(M)})$, and the definition of the probability vectors $p_{\epsilon}(x)$ follows from (10). The split-step scheme (21) generalises along the same lines.

### v3

@@ 3. Approximating the conditional mean
In order to implement (5), we need to define $m(x;\epsilon)$ and $\Sigma(x)$ for any $x\in\mathbb{R}^{d}$. In this section, we discuss how one can obtain these functions from the training samples $\{x^{(i)}\}_{i=1}^{M}$ and the Markov chain approximation (10).
We introduce the vector-valued function $t(x)\in\mathbb{R}^{M}$ with entries
(14) $t_{i}(x)=\exp\left(-\frac{1}{2\epsilon}(x^{(i)}-x)^{\top}\left(K(x)+K(x^{(i)})\right)^{-1}(x^{(i)}-x)\right)$
for $i=1,\ldots,M$. We then define the probability vector using the Sinkhorn weights, $v$, obtained in (10), i.e.,
(15) $p(x)=\frac{D(v)t(x)}{v^{\top}t(x)}\in\mathbb{R}^{M}$
for all $x\in\mathbb{R}^{d}$. This vector gives the transition probabilities from $x$ to training samples $\{x^{(i)}\}_{i=1}^{M}$ and provides a finite-dimensional approximation to the conditional probability distribution $\pi_{\epsilon}(\cdot|x)$ of the true underlying diffusion process; i.e., the semi-group $\exp(\epsilon\mathcal{L})$ with generator $\mathcal{L}$ corresponding to the generalized reversible diffusion process (7).
Finally, using the probability vector $p(x)$, which implicitly depends on $\epsilon$, and introducing the data matrix of training samples
@@
The construction of the conditional mean $m(x;\epsilon)$ is known as the barycentric projection of the entropy-optimally coupling [50, 44]. In optimal transport, $v$ plays the role of the optimizer of the dual problem. We also mention the connection to de-noising schemes which utilize a structure similar to (17). See the recent review [40].
@@ 3.1. Sampling algorithms
We now present our main Langevin sampling strategies based on the previously introduced probability vector $p(x)$ in (15). Sampling schemes of the form (5) have the same drift term (17), but differ in the way the diffusion matrix $\Sigma(x)$ is defined. We consider a data-unaware diffusion as well as a data-aware diffusion which turns out to be advantageous in generating new samples from the data distribution $\pi$ (see the numerical experiments in Section 5).
@@ 3.1.1. Langevin sampler with data-unaware diffusion
(18) $X_{n+1}=X_{n}+\Delta\tau\left(\frac{m(X_{n};\epsilon)-X_{n}}{\epsilon}\right)+\sqrt{2K(X_{n})}\Xi_{n}$
@@
(19) $s(x)=\frac{m(x;\epsilon)-x}{\epsilon}$
@@
Let us briefly discuss the qualitative behavior of the time-stepping method (20) as a function of $\epsilon>0$. For large $\epsilon$, the expected value $m(x;\epsilon)$ will become essentially independent of the current state $X_{n}$ and the diffusion process will sample from a centred Gaussian. For $\epsilon\to 0$, on the other hand, the probability vector $p(x)$ can potentially degenerate into a vector with a single entry approaching one with all other entries essentially becoming zero. Hence a key algorithmic challenge is to find a good value for $\epsilon$ and a suitable $K(x)$, which guarantee both good mixing and accuracy, i.e., $X_{n}\sim\pi$ as $n\to\infty$.
In terms of initialisation, it is often best to initialise from one of the training data $x^{(j^{\ast})}$ with $j^{\ast}\in\{1,\ldots,M\}$ chosen at random.
@@ 3.1.2. Langevin sampler with data-aware diffusion
(21) $C(x)=\epsilon^{-1}(\mathcal{X}-m(x;\epsilon)1_{M}^{\top})D(p(x))(\mathcal{X}-m(x;\epsilon)1_{M}^{\top})^{\top}\in\mathbb{R}^{d\times d},$
which is the (scaled) covariance matrix associated with the probability vector $p(x)$. Here $1_{M}\in\mathbb{R}^{M}$ denotes the $M$-dimensional vector of ones. Therefore, one can more directly implement a Gaussian approximation associated with the transition probabilities $p(x)$ and introduce the update
(22) $X_{n+1}=X_{n}+\Delta\tau\left(\frac{m(X_{n};\epsilon)-X_{n}}{\epsilon}\right)+\sqrt{C(X_{n})}\Xi_{n}.$
@@
(23) $X_{n+1}=m(X_{n};\epsilon)+\sqrt{C(X_{n})}\Xi_{n},$
which we found to work rather well in our numerical experiments since it directly captures the uncertainty contained in the data-driven coupling $P$. The scheme (23) corresponds to setting $\Sigma(x)=C(x)$ in (5). Also note that the scheme (23) still depends on $K(x)$ through the probability vector $p(x)$.
@@ 3.2. Algorithmic properties
We briefly discuss several important properties on the stability and the ergodicity of the proposed Langevin samplers. The following Lemma establishes that, since each $p$ is a probability vector, $m(x;\epsilon)=\mathcal{X}p(x)$ is a convex combination of the training sample $\{x^{(i)}\}_{i=1}^{M}$.
@@
The Lemma follows from the definition (17), which we write as $m(x;\epsilon)=\sum_{j=1}^{M}x^{(j)}p_{j}(x)$, and the fact that $p(x)$ is a probability vector with $0\leq p_{j}(x)\leq 1$ for all $\epsilon>0$ and all $x\in\mathbb{R}^{d}$. ∎
@@
Let us assume that the data generating density $\pi(x)$ has compact support. Then the time-stepping method (20) possesses a unique invariant measure and is geometrically ergodic provided the norm of the symmetric positive matrix $K(x)$ is bounded from above and below for all $x\in\mathbb{R}^{d}$.
@@
We consider $K(x)=I$ for simplicity. We introduce the Lyapunov function $V(x)=\|x\|^{2}$ and balls
(25) $\mathcal{B}_{R}=\{x\in\mathbb{R}^{d}:\|x\|\leq R\}$
of radius $R>0$ in $\mathbb{R}^{d}$. Since $m_{\epsilon}(X_{n})\in\mathcal{C}_{M}$ and $\pi(x)$ has compact support, one can find radii $R_{\ast}>0$ and $R>\sqrt{R_{\ast}^{2}+2\epsilon}$, which are independent of the training data $\{x^{(i)}\}_{i=1}^{M}$, such that $\mathcal{C}_{M}\subset\mathcal{B}_{R_{\ast}}$ and
(26) $\mathbb{E}[V(X_{n+1})|X_{n}]\leq\lambda V(X_{n})$
for all $X_{n}\notin\mathcal{B}_{R}$ and suitable $10\leq\lambda<$ 1. This follows from the fact that $\mathbb{E}[V(X_{n+1})|X_{n}]<R_{\ast}^{2}+2\epsilon$, while $V(x)>R^{2}$ for all $x\notin\mathcal{B}_{R}$. One then chooses $\lambda=(R_{\ast}^{2}+2\epsilon)/R^{2}<1$. Furthermore, there is a constant $\delta>0$ such that
${\rm n}(x^{\prime};m(x;\epsilon),2\epsilon I)\geq\delta$
for all $x,x^{\prime}\in\mathcal{B}_{R}$. Here ${\rm n}(x;m,\Sigma)$ denotes the Gaussian probability density function with mean $m$ and covariance matrix $\Sigma$. In other words, $\mathcal{B}_{R}$ is a small set in the sense of [39]. Geometric ergodicity follows from Theorem 15.0.1 in [39]. See also the self-contained presentation in [37]. ∎
@@
Lemma 3.3.
If $K(x)=I$, the conditional mean estimator $m(x;\epsilon)$ is equivalent to
(27) $m(x;\epsilon)=x+\epsilon\nabla_{x}\log\Pi(x;\epsilon),\qquad\Pi(x;\epsilon):=(v^{\rm T}t(x))^{2}.$
Hence, the update (20) is equivalent to an Euler–Maruyama discretization of Langevin dynamics (1) with modified probability density
(28) $\tilde{\pi}(x;\epsilon):=\frac{\Pi(x;\epsilon)}{\int\Pi(x;\epsilon)\,{\rm d}x}.$
In other words, for $\epsilon$ sufficiently small and $K(x)=I$, (20) samples approximately from $\tilde{\pi}(x;\epsilon)$.
Proof.
Formula (27) follows from (15) and (14). ∎
@@
(30b) $X_{n+1}$ $=m(X_{n+1/2};\epsilon).$
These split-step schemes have been used in our numerical experiments. We note that (29b) can be viewed as a de-noising step applied to a noisy $X_{n+1/2}$. See [40] for a recent survey of de-noising techniques used in image processing and their connection to SGM.
Figure 1. Histograms of the $x_{1}$- and $x_{2}$-components of the training as well as generated data for Example 3.1. One finds that the split-step scheme effectively denoises the $x_{2}$-component while faithfully reproducing the standard normal distribution in the $x_{1}$-component.
Example 3.1.
We return to Example 1.1 and demonstrate the performance of the proposed split-step scheme (29) for this simple linear problem. We sample $M=10^{3}$ training samples from the Gaussian distribution with mean zero and covariance matrix
(31) $C=\left(\begin{array}[]{cc}1+\nu&0\\ 0&\nu\end{array}\right)$
with $\nu=10^{-4}$. The split-step scheme (29) is implemented with $\epsilon=0.1$ and a total of $10^{5}$ samples are generated. See Figure 1 for numerical results. It is found that the proposed scheme samples correctly from the normal distribution in $x_{1}$ and, at the same time, denoises the $x_{2}$-component. We recall that a standard Euler–Maruyama discretisation would require step-sizes $\Delta\tau<0.0002$ and would sample from the noisy distribution ${\rm N}(0,C)$.
We finally discuss the computational complexity of the proposed algorithms in terms of the size, $M$, of the training data and the dimension, $d$, of the samples space $\mathbb{R}^{d}$. In the off-line phase of the algorithm, the accelerated Sinkhorn algorithm of [58] converges essentially in order $\mathcal{O}(M^{0})$ iterations and each iteration requires matrix-vector multiplications; hence being of order $\mathcal{O}(M^{2})$. Computing $T$ involves calculating $\mathcal{O}(M^{2})$ distances between vectors in $\mathbb{R}^{d}$. The online phase requires the multiplication of the $d\times M$-dimensional data matrix $\mathcal{X}$ with the $M$-dimensional weight vector $p(x)$. The computation of $p(x)$ in turn involves the computation of $M$ distances $\|x-x^{(j)}\|^{2}$ and inner products of $M$-dimensional vectors. Overall the computational complexity of the online phase is of order $\mathcal{O}(dM)$.
The more problematic part is the accuracy of the Schrödinger bridge approximation to the semi-group $\exp(\epsilon\mathcal{L})$. Optimal scaling of $\epsilon$ as a function of $M$ leads to an approximation error of order $\mathcal{O}(M^{-2/(8+d)})$, which requires an exponential increase of the sample size $M$ as a function of the dimension, $d$, of sample space in order to reach a desired accuracy. While improvements in accuracy can be obtained by the variable bandwidth techniques discussed next for low dimensional problems, a localized formulation of the Schrödinger bridge sampler, as proposed in [22], can lead to an effectively dimension-independent accuracy.
@@ 3.3. Variable bandwidth diffusion
It is an active area of research to select a $\rho$ that increases the spectral gap of the associated generator $\mathcal{L}$, given by
(33) $\mathcal{L}f=\pi^{-1}\nabla\cdot(\pi\rho\nabla f)=\nabla\cdot(\rho\nabla f)+\rho\nabla\log\pi\cdot\nabla f,$
while not increasing computational complexity. A larger spectral gap implies a faster convergence rate [46], indicating that the generated samples are closer to the reference at a finite time, exhibiting a high accuracy. We demonstrate numerically in Section 5 that $\rho$ can indeed be used to increase the sampling accuracy. More specifically, the bandwidth $\rho(x)$ is chosen as
@@
(35) $\tilde{\pi}_{i}=Z^{-1}\pi_{i},\qquad Z:=\frac{1}{M}\sum_{j}\pi_{j}.$
@@
(36) $\rho_{i}=\tilde{\pi}_{i}^{\beta}=Z^{-\beta}\pi_{i}^{\beta}$
@@
In order to derive an appropriate time-stepping scheme, we note that the drift term in the Itô formulation (7) with $K(x)=\rho(x)I$ can be expressed as
@@
and, hence, it holds that
(39) $m(x;\epsilon)=\mathcal{X}p(x)\approx\exp(\epsilon\mathcal{L})\,\mbox{id}(x)=\mu(x;\epsilon),$
and
(40) $\exp(\epsilon\mathcal{L})\,\mbox{id}(x)\approx x+\epsilon\mathcal{L}\,\mbox{id}(x)=x+\epsilon s(x)$
as desired (cf. (19)). We emphasize that the drift now satisfies $s(x)=\rho(x)\nabla\log\pi(x)+\nabla\rho(x)$ and is no longer equivalent to the score $\nabla\log\pi(x)$. For the numerical variable bandwidth experiments conducted in Section 5, we therefore used (29) with $K(x)=\rho(x)I$. We remark that our aim is not to reproduce the actual dynamics of an underlying stochastic process, i.e. the process associated with the generator (33), but rather to sample from the distribution $\pi(x)$.
@@ 4. Bayesian inference and conditional sampling
where $V(x)$ denotes a known potential such as the negative log-likelihood function in case of Bayesian inference with $\pi(x)$ taking the role of the prior distribution. We obtain, for example, the adjusted time-stepping scheme
(42) $X_{n+1/2}$ $=X_{n}-\epsilon\nabla V(X_{n})+\sqrt{2}\,\Xi_{n},$
(43) $X_{n+1}$ $=m(X_{n+1/2};\epsilon)$
@@
(44) $x^{\ast}=\arg\min_{x}\left\{V(x)-\log\pi(x)\right\}$
for given potential $V(x)$ and regulariser $-\log\pi(x)$. Again assuming that only samples $\{x^{(i)}\}_{i=1}^{M}$ of $\pi(x)$ are available, an algorithm for approximating $x^{\ast}$ can be defined as follows:
(45a) $x_{n+1/2}$ $=x_{n}-\epsilon\nabla V(x_{n}),$
(45b) $x_{n+1}$ $=m(x_{n+1/2};\epsilon).$
Here $m(x;\epsilon)=\mathcal{X}p(x)$ approximates the optimization update associated with the regulariser $-\log\pi(x)$. The iteration is stable and any limiting point $x_{\infty}$ is contained in the convex hull of the data points $\{x^{(i)}\}_{i=1}^{M}$.
The sampling schemes (29) and (30) can furthermore be extended to conditional generative modeling. More specifically, consider a random variable $x=(y,z)$, which we condition on the first component for given $y=y^{\ast}$. We assume here that the decomposition of $x$ into the parts $z$ and $y$ is given, and we wish to sample from $\pi(z|y^{\ast})$ given samples $x^{(i)}=(y^{(i)},z^{(i)})$, $i=1,\ldots,M$, from the joint distribution $\pi(x)=\pi(y,z)$.
In order to perform the required conditional sampling, we propose a method which combines approximate Bayesian computation (ABC) with our diffusion map based sampling algorithm. As before, we construct vectors of conditional expectation value $m(x;\epsilon)$ based on the samples $\{x^{(i)}\}_{i=1}^{M}$. We assume that the bandwidth parameter $\epsilon$ used in the diffusion map approximation is also applied in the ABC misfit function, i.e.,
(46) $L(y,y^{\ast})=\frac{1}{2\epsilon}\|y-y^{\ast}\|^{2}.$
This suggests the following split-step approximation scheme. Given the last sample $X_{n}=(Y_{n},Z_{n})$, we first update the $y$-component using
(47) $\hat{Y}_{n}=Y_{n}-\epsilon\nabla_{y}L(Y_{n},y^{\ast})=y^{\ast}.$
In other words, we replace the current $X_{n}=(Y_{n},Z_{n})$ by $\hat{X}_{n}=(y^{\ast},Z_{n})$. Next we apply the split-step scheme (29) to $\hat{X}_{n}$, i.e.,
(48) $X_{n+1/2}$ $=\hat{X}_{n}+\sqrt{2K(\hat{X}_{n})}\Xi_{n},$
(49) $X_{n+1}$ $=m(X_{n+1/2};\epsilon),$
where $m(x;\epsilon)=\mathcal{X}p(x)$ with $\mathcal{X}=(x^{(1)},\ldots,x^{(M)})$, and the definition of the probability vectors $p(x)$ follows from (15). The split-step scheme (30) generalises along the same lines.
We show in Section 5.3 how conditional sampling with the Schrödinger bridge enables drawing inaccessible random variables compatible with a known macroscopic state, a problem known as stochastic subgrid-scale parametrization. We further employ conditional sampling in Section 5.4 to draw random trajectories of a dynamical system where a future state is conditioned on the current state.